Please help with these questions. 2. If the position vector r= 151 - 10j +20k (m) and force F = 250i +400j - 600k (N). a) Find the angle between r and F in deg
Answer #1

2 q=15a logt + 20 i (m) 7 = 250 î+400ģ- Gook (N) Let angle between 7 and F be d. Å F = 1 All Flead [By det product Al 15²+10²6 2 * I! 21 // 3 fexs ch Total area of the shaded fertion is A = D + (8 x 6) Fixers] <quaverian = 48.86 sq. in. Let distanIM Ix xixo 10x1x (8.5-5.412)² +(4x846) + 8x6x (5.412-4) 2 x (5.412-4) + in >> 4 (26.19 +351.699 -7.049] in = 440.34 in 4 ::